### v1

# miniCTX: Neural Theorem Proving with

Abstract
We introduce miniCTX, which tests a model’s ability to prove formal mathematical theorems that depend on new definitions, lemmas, or other contextual information that was not observed during training. miniCTX contains theorems sourced from real Lean projects and textbooks, each associated with a context that can span tens of thousands of tokens. Models are tasked with proving a theorem given access to code from the theorem’s repository, which contains context that is helpful or needed for the proof. As a baseline for miniCTX, we introduce file-tuning, a simple recipe that trains a model to generate a proof step conditioned on the preceding file contents. File-tuning substantially outperforms the traditional neural theorem proving approach that fine-tunes on states alone. Additionally, our file-tuned model improves performance on the standard miniF2F benchmark, achieving a pass rate of 33.61%, which is a new state-of-the-art for 1.3B parameter models. Alongside miniCTX, we offer ntp-toolkit for automatically extracting and annotating theorem proving data, making it easy to add new projects into miniCTX to ensure that contexts are not seen during training. miniCTX offers a challenging and realistic perspective on evaluating neural theorem provers. [^1] Data and models: HuggingFace. ntp-toolkit: cmu-l3/ntp-toolkit. Evaluation code: cmu-l3/minictx-eval.
1 Introduction
Formal theorem proving in interactive theorem provers (ITPs) provides a testbed for evaluating the reasoning capabilities of large language models (LLMs). Theorem proving capabilities can then directly translate to automation for mathematicians, such as via tools that complete or formalize proofs [1, 2, 3, 4]. However, despite their promise, we see a gap between current language model-based provers and the complexity of real-world theorem proving.
Our motivating observation is that theorems and proofs depend on various forms of context, such as newly-defined definitions and lemmas. For instance, to prove results about a square, one might first formalize a definition of a rectangle, prove some results about rectangles, then specialize them to a newly-defined square [5] (Figure 1). However, existing methods for training and evaluating LLM-based theorem provers do not take such context into account. For example, benchmarks often focus on proving standalone competition problems (e.g., miniF2F [6]) or theorems from a library that the model has trained on (e.g., mathlib [7, 8]), and LLM-based provers are trained to accept only a proof state as input, making them unaware of new theorems and definitions [9].
From a language modeling perspective, a key challenge is the sheer size of potentially useful context. For instance, proving a theorem in the Prime Number Theorem project without having seen it during training can require reasoning over a long context of interdependent definitions and theorems [10]. Therefore, theorem proving with newly-defined contexts offers a challenging testbed for long-context reasoning, in addition to more accurately reflecting practical proving. With these considerations in mind, we define a new setting of context-dependent theorem proving, and develop an evaluation framework, benchmark, and training recipe for theorem proving with context.
2 Theorem proving with context
Formal theorem proving involves two stages: defining mathematical objects and facts relevant to the desired result, then stating and proving the result itself. For language model-based provers to function as useful tools in this real-world setting, they need to be able to work with new information such as new definitions or lemmas. For example, a system suggesting proofs in the Prime Number Theorem project should be familiar with the project’s definition of “square”. Unfortunately, current language model-based provers are trained on a static snapshot of data, and are hence unaware of any context that was created after the model was trained (Figure 1). As a result, it remains unclear whether these models can work with new information, and how to enable this capability.
Figure 1: Traditional language model-based provers train on a static collection of data (e.g., a snapshot of Mathlib), and are evaluated on standalone theorems that depend only on facts and theorems seen during training (e.g., competition problems [6] that rely only on basic facts in algebra and number theory). We argue that neural theorem provers must also operate in the realistic context-dependent setting, in which results depend on working with new mathematical objects and their facts (right).
Context-dependent proving.
We study context-dependent theorem proving, where the goal is for a model to generate proofs $y$ for new theorems $x$, based on a context $c$ that includes background information such as definitions, lemmas, or natural language comments. Formally,
$\arg\max_{M}\mathbb{E}_{(x,c)\sim p}\mathbb{E}_{y\sim M(\cdot|x,c)}v(x,c,y),$ (1)
where $(x,c)\sim p$ denotes a (theorem, context) pair from a context distribution $p$, $M$ is a model that produces a proof $y$, and $v$ returns 1 if the proof is correct and 0 otherwise (implemented by Lean [^2] We focus on the Lean [11] proof assistant in this paper, though the problem of context-dependent proving is present in other proof assistants such as Coq [12] and Isabelle [13].).
In this paper, we treat a Lean repository as a context distribution. Each context $c$ is some subset of code in the repository that does not contain the theorem’s formal proof. In this paper we take $c$ to be the code that precedes the theorem $x$ in a file. This context includes new definitions, lemmas, and comments that are relevant to the theorem but does not show the theorem’s formal proof.
Evaluating context-dependent proving.
To measure a model’s performance on context-dependent proving, a model is tasked with generating a proof $y$ given a theorem $x$ and its preceding code $c$. Given a language model, we can test three kinds of generalization by ensuring the following:
  • Theorem-level generalization: the proof must not occur in the model’s training data.
  • Context-level generalization: the code $c$ and proof must not occur in the training data.
  • Project-level generalization: the entire repository must not occur in the training data.
Assuming that we know the date at which a model was trained, we can guarantee theorem-level, context-level, and project-level generalization by selecting theorems, files, or projects (respectively) that were added after the model’s training date.
To realize our evaluation framework, we develop a method that automatically extracts a dataset of (context, theorem, metadata) examples from a Lean repository, i.e., $f(R)\rightarrow\{(c,x,z)_{i}\}_{i=1}^{N}$. Next, we apply this method to create miniCTX, a benchmark for context-dependent theorem proving.
3 miniCTX: a benchmark for theorem proving with context
We develop miniCTX, a Lean 4 theorem proving benchmark of theorems that depend on newly-defined lemmas, definitions, and proofs from within a project. miniCTX is currently based on 376 theorems from four projects: recent Lean formalizations of the Prime Number Theorem (PrimeNumberTheoremAnd[10] and the Polynomial Freman-Ruzsa Conjecture (PFR[14], problems from an introductory text on theorem proving (HTPI[15], and recent results from Lean’s mathematical library (Mathlib[16].
Each theorem in miniCTX is accompanied by the following data, formatted in JSON:
  1. 1.
    Theorem statement,
  2. 2.
    Preceding file contents up to the theorem statement,
  3. 3.
    Metadata, including:
    1. (a)
      File name,
    2. (b)
      Project commit and version,
    3. (c)
      Commit at which the theorem and its file was added,
    4. (d)
      Position of the theorem and number of premises preceding it,
    5. (e)
      Proof length and type,
    6. (f)
      Whether the statement or proof uses definitions or lemmas from within the file or repository.
See example in §A.1. We make the dataset available at l3lab/miniCTX, along with corresponding evaluation code based on the Lean REPL [17]. The dataset has an Apache 2.0 license.
Prime Number Theorem.
PrimeNumberTheoremAnd [10] is a project started in January 2024 that formalizes the prime number theorem in Lean as well as related concepts, such as residue calculus on rectangles in $\mathbb{C}$. We find the files Rectangle.lean and ResidueCalcOnRectangles.lean suitable for our purpose of testing context-dependent theorem proving, especially when we use preceding file content as context, as each file is self-contained within the project and contains new definitions (rectangles, squares) and many interdependent lemmas. See §A.2 for an illustration of such lemmas. In addition, most theorems from ResidueCalcOnRectangles.lean rely on the definitions from Rectangle.lean, which serves as a perfect example of cross-file dependencies. We extracted 87 theorems from these files. Assuming that a model was trained prior to January 2024, this split guarantees the evaluation of project-level, context-level, and theorem-level generalization.
PFR.
PFR [14] is a project started in November 2023 that formalizes a proof of the Polynomial Freiman–Ruzsa (PFR) conjecture. We included 54 theorems from PFR. We find that proofs of theorems in PFR tend to be much more monolithic and longer in length than those in Mathlib or other libraries. PFR also defines custom mathematical concepts and notations (such as Ruzsa distance) and a proof typically depends on many lemmas in PFR outside the current file. All of the theorems were added to PFR after November 2023. Assuming that the model was trained prior to this date, this split guarantees the evaluation of project-level, context-level, and theorem-level generalization.
Recent Mathlib Commits.
Lean’s mathematical library, Mathlib [16], is a community-maintained Lean repository including mathematical concepts as well as programming APIs and common tactics. It is the single largest Lean library that users contribute to, and is therefore representative of the production environment in which neural theorem provers are deployed. Mathlib is a long-standing project, and it is common practice to train language model-based provers on Mathlib. It is therefore likely that Mathlib source files have been observed during training. However, Mathlib is frequently updated, with new definitions, theorems, and refactorings occurring on a daily basis. Hence we can test theorem-level generalization by tasking a model with proving newly added theorems, given a context that may or may not have been observed during training.
We included 50 theorems added to Mathlib in April 2024, by filtering recent Mathlib commits to ones that only add new theorems. Many of the theorems added are simple lemmas that depend on earlier ones (e.g., ones seen during training). As Mathlib generally refactors new theorems by breaking down long theorems to shorter lemmas, most new theorems are not difficult to prove, and give a realistic representation of where neural theorem provers are used. Assuming that the model was trained prior to April 2024, the Mathlib split guarantees the evaluation of theorem-level generalization.
HTPI.
HTPI contains the Lean code for the book How to Prove It (HTPI) [18], which explains a systematic approach to constructing mathematical proofs with Lean. It covers various topics, including elementary logic, number theory, and proving techniques like mathematical induction, along with their implementation in Lean. As supplementary material to the textbook, the problems in HTPI are formulated in a similar fashion: the files typically start with basic definitions and lemmas that might be used throughout the entire file, followed by exercises and several example problems. [^3] The main chapter and exercises are separated in the original project: HTPILeanPackage. We manually merged them for evaluation in the fork: HTPILeanPackage4.7. Therefore, models can utilize definitions, lemmas, and proof structures from example problems to solve similar exercises. Intuitively, the model must understand and apply the context provided within each file, making it an effective benchmark for testing context-aware theorem-proving models.
Split Problems Avg. Context Length (tokens) Avg. Proof Steps
miniF2F [[6]] Valid/Test 488     153*    3.0**
miniCTX Prime 87   10,630    3.6
PFR 54   17,495 27.7
Mathlib 50   14,440    6.1
HTPI 185   39,050    10.7**
All 376   26,106 10.9
* Average context length for MiniF2F only counts libraries and definitions.
** Average proof steps for MiniF2F and HTPI only include problems with proofs.
Table 1: Overview of Problem Statistics in miniF2F and miniCTX.
3.1 Core challenges: dependencies, length, constraints, and varied content
miniCTX introduces several challenges for language-model based provers.
File and repository dependencies. First, a theorem’s statement or proof may have contextual dependencies, i.e., requiring the use of definitions or lemmas from the new repository. Assuming that the model has not observed the repository during training, a model must have some mechanism for taking in the new definitions and lemmas, and must have the ability to produce proofs that, intuitively, require understanding the new context. As a proxy indicator of having dependencies, we record whether the statement or proof uses definitions or theorems from within the file or repository.
Long context length.
A second challenge is the length of the context in miniCTX. As seen in Table 1, the average context length is over 26,000 tokens. Intuitively, the model must reason over long contexts containing information that is needed to produce a proof. This differs from proving standalone competition problems in miniF2F, which have an average context length of 153 tokens.
Constraints. Third, real-world proofs may have particular constraints, such as disabling common automation tools and defining new tactics, or custom definitions and syntax such as in PFR. A model must be able to adapt and produce these context-dependent proofs.
Varied content and difficulty. A final challenge is that miniCTX has varied mathematical content in its theorems, thus requiring a flexible automated proving system. This ranges from writing explicit calculation proofs in HTPI to writing sophisticated proofs in PFR. One axis of variation is the length of the human-written proof (if available), which we record and treat as a proxy for difficulty.
Helpful context. We also note that context may make some theorems easier to prove compared to standalone theorems such as competition problems. For instance, Lean source code can have natural language comments or Latex proofs that may help constrain the space of possible proofs. Moreover, some proofs within a repository or file often have analogous patterns or structure, which intuitively may make subsequent theorems easier to prove (e.g., see §A.2).
Automatically updating the benchmark. miniCTX’s format is amenable to periodically updating the benchmark with new projects to ensure that theorems, definitions, or full projects are not seen by language models trained prior to a particular date. Future updates will be automatically extracted from new Lean projects and commits using ntp-toolkit (§4.1). Users can also create their own evaluation data with ntp-toolkit. To test a model’s ability to adapt to in-file contexts, we recommend using files that are long, but do not contain intra-repository imports. For example, Rectangle.lean from the Prime Number Theorem contained 47 theorems and only imported Mathlib packages.
4 File-tuning for theorem proving with context
Writing a proof can be seen as a sequential process $(x_{1},y_{1}),(x_{2},y_{2}),\ldots$ of states $x_{t}$ and tactics $y_{t}$. A state contains what is left to prove (the goal), and available information (the hypotheses). A tactic transitions the proof to a new state. If the state contains no remaining goals, the proof is complete. Concretely, a user applies tactics by writing Lean code, Lean keeps track of the state, and the development environment shows the state and the written code.
Figure 2: State-tactic vs. file tuning.
The traditional approach to training a language model for theorem proving is to train a model on (state, tactic) pairs, i.e., train it to predict the next step of a proof (i.e., the tactic), given the state provided by the proof assistant (i.e., the proof state) [9, 7, 19, 8]. A drawback to this approach is that at test time, the model is not aware of new context outside of the proof state, such as new lemmas. We will see later on that models trained with this state-tactic approach fail at context-dependent proving.
As a stronger baseline for context-dependent proving, we present file-tuning, a simple recipe that trains with (preceding file context, proof state, next-tactic) tuples instead of training with (proof state, next-tactic) pairs (Figure 2). This lets the model use new definitions, theorems, or other information that are defined prior to the current tactic at training or at test time. In practice, file-tuning requires extracting contexts and proof states from Lean. We describe our toolkit for doing so next.
4.1 ntp-toolkit for automated data extraction and evaluation
We provide ntp-toolkit to extract training data, evaluation data, and interactive Lean evaluation.
Data extraction.
ntp-toolkit contains a general-purpose data extraction tool that extracts examples from an arbitrary Lean 4 repository and formats them into examples for language-model fine-tuning. The tool is implemented in Lean based on Kim Morrison’s lean-training-data [20]. Specifically, ntp-toolkit takes in a configuration file with one or more Lean repositories specified. Each repository is transformed into next-tactic and full proof examples stored in JSON Lines files. The next-tactic data is suitable for making file-tuning examples of the form (context, state, next-tactic). The full proof data is suitable for making evaluation or training examples of the form (context, theorem, proof). See Appendix §C.1 for the JSON formats provided by the data extraction.
Instruction-tuning data.
ntp-toolkit re-formats extracted data into input-output examples for file-tuning and state-tactic tuning. Each example is preceded with a natural language instruction that describes the task (generating a next-tactic) and the available inputs (e.g., file contents, proof state). Data of this form is often referred to as “instruction-tuning data”. The specific instructions and formats that we used are shown in the Appendix (§C.2). Users can also adjust the formats as needed.
Interaction via Lean REPL.
Proving a theorem with a language model requires interacting with Lean: sending tactics, receiving error messages, etc. In an effort to better integrate LLMs and Lean, we developed a Python wrapper for the Lean Read-Eval-Print Loop (REPL) [17]. The Lean REPL offers an interactive environment for submitting code to and receiving messages from a Lean instance. We provide a Python interface for (1) submitting both complete Lean files and individual tactics, and (2) receiving and handling feedback from Lean, such as error messages or indicators that a tactic application is valid or a proof is complete. This interface enables standard tactic-based proving along with other strategies (e.g., submitting and checking a full proof).
5 Experiments
First, we describe our procedure for file-tuning a language model on examples extracted from Mathlib. Second, we evaluate the file-tuned model and other baselines on miniCTX. We study performance along various axes, such as dependencies, context length, difficulty, and the type of information in the context. Our experiments show that traditional state-tactic models struggle in the context-dependent setting compared to file-tuned models that can use context at proving time. Finally, our investigation reveals several open challenges and future directions that we discuss in §5.5
5.1 File-tuning on Mathlib
Data extraction.
We ran ntp-toolkit’s next-tactic extraction on a snapshot of Mathlib, yielding 307,049 examples available at l3lab/ntp-mathlib. We then ran ntp-toolkit’s instruction tuning script on these examples, yielding file-tuning examples and state-tactic examples. For the file-tuning examples, as an initial method for handling the long Lean files, we either truncate the middle of an input file so that the file contents is 1024 tokens, or take only the preceding 1024 tokens, with the strategy selected at random for each example. The state-tactic examples are at l3lab/ntp-mathlib-instruct-st. The union of file-tuning and state-tactic examples are at l3lab/ntp-mathlib-instruct-context, split into 583k train, 15k dev, and 15k test examples.
File-tuning.
Next, we fine-tune a language model on the union of file-tuning and state-tactic examples. We use the DeepSeek-Coder-1.3b language model [21] based on its performance on code generation tasks and its size, which allowed us to fine-tune with our computational budget. We fine-tune the model for 3 epochs and select the model based on held-out validation perplexity evaluated every 4,000 steps. The model is available at l3lab/ntp-mathlib-context-deepseek-coder-1.3b.
State-tactic tuning. We fine-tune a similar model on l3lab/ntp-mathlib-instruct-st using the same hyperparameters. The model is available at l3lab/ntp-mathlib-st-deepseek-coder-1.3b.
5.2 Evaluation setup
We evaluate models for the task of tactic-based theorem proving: given a theorem statement, a model generates one tactic at a time while receiving states from the proof assistant. We use a standard best-first search strategy [9, 7, 22, 8, 1] which prioritizes partial proofs based on the model’s average log probabilities. This search method is parameterized by the number of generated tactics per iteration $S$, and the maximum number of iterations $T$. We use the setting from [23, 1] ($S=32$, and $T=100$).
We evaluate five types of baselines: (1) pass@1 full proof generation using GPT-4o: we prompt the model with only the theorem statement and require it to generate a complete proof (see Appendix (§C.2) for details of the prompts and few-shot examples); (2) pass@1 full proof generation with in-file context using GPT-4o: we supplement the theorem statement with up to 8000 tokens of in-file context; (3) the file-tuning model described in (§5.1); (4) the state-tactic model described in (§5.1); and (5) a state-tactic prompting model: we prompt a pre-trained language model with (state, tactic) examples, following [23]. We use the Llemma-7b model [23].
5.3 Top-level results
File-tuning maintains performance on competition problems.
First, we evaluate on miniF2F, a standard benchmark based on competition problems that do not require the use of context. We adapt it to the context-dependent setting by simply providing the model with the miniF2F import statements. As seen in Table 2, the state-tactic models achieve 32.79%, which is higher than existing methods tuned on LeanDojo [8] (e.g., 26.50% in  [8], 27.90% in  [1]). This validates the efficacy of ntp-toolkit’s data extraction and our fine-tuning. The file-tuned model achieves 33.61%, showing that file-tuning can maintain or even improve performance in a setting that does not require context.
File-tuning substantially improves context-dependent proving.
Second, Table 2 shows results on miniCTX. These proofs typically depend on more context than those in miniF2F. We see a dramatic improvement for the file-tuned model over the state-tactic methods. The file-tuned model achieves 31.65% compared to 9.31% from the state-tactic model. The gap shows the importance of evaluating the context-dependent setting and theorem proving methods that use context beyond the proof state.
MiniF2F MiniCTX
Models Test Prime PFR Mathlib HTPI Avg.
GPT-4o (full proof) - 1.15% 5.56% 2.00% 9.73% 5.59%
GPT-4o (+ context) - 13.79% 1.85% 18.00% 31.89% 22.07%
State-tactic prompting 28.28% 19.54% 5.56% 16.00% 19.15% 20.61%
State-tactic tuning 32.79% 11.49% 5.56% 22.00% 5.95% 9.31%
File tuning 33.61% 32.18% 5.56% 34.00% 38.38% 31.65%
Table 2: Performance comparison of different models on MiniF2F and MiniCTX.
Figure 3: Performance partitioned by dependency type. File-tuned models substantially outperform state-tactic tuned models on theorems with definition and/or theorem dependencies.
5.4 Analysis
We analyze the file-tuned and state-tactic models on miniCTX further along several axes, including the kinds of contextual dependencies, the difficulty, and the content made available in the context.
File-tuning especially helps on problems with dependencies.
We use the miniCTX metadata to categorize theorems from Prime, PFR and Mathlib splits based on their in-file dependencies. Figure 3 shows the performance on problems that depend on in-file definitions, in-file theorems, both, or neither. We also show miniF2F as an additional reference point for problems without in-file dependencies. The file-tuned model shows a marked improvement over the state-tactic tuned model in all scenarios that have dependencies (outlined in black). The gap is much larger than on problems without dependencies. We conclude that file-tuning specifically helps in the realistic setting of proving with new definitions and theorems.
Some definition dependencies can be handled given access only to the proof state.
Interestingly, the state-tactic tuned model performs decently on definition-only problems, achieving a success rate of 27.66% despite lacking direct access to explicit definitions. Upon closer inspection, we recalled that the goal state provided by Lean includes the names and types of definitions, allowing the model to call automation tactics such as simp with correct arguments (See Appendix §D.5).
Context Type Accuracy (%)
No Context 17.02
Imports & Definition 29.79
Theorems 38.29
No proof 46.81
All 46.81
Table 3: Ablating components of the context.
Definitions and theorems in the context are both important.
To determine the relative importance of various context components, in Table 3 we evaluated the file-tuned model’s performance under two modified settings: one where only in-file definitions are provided, and another where only in-file theorems are included. Performance drops when the model is restricted to using either component in isolation compared to the full context, which includes both definitions and theorems.
Suprisingly, pass rate was similar without proofs in the context.
We also investigate the impact of proofs of previous theorems in the context. We modified the theorem proving context of Rectangle.lean file by replacing all proofs associated with theorems in the context with the placeholder “sorry”, while retaining the theorem statements. Surprisingly, the pass rate remained the same as the original setting in which full proofs were available (though, the models have slightly different sets of successful problems) However, we inspected some proofs and found that our model did have some ability to replicate the proof strategies outlined in the context, as illustrated in Appendix §D.3.
File-tuning improves across all difficulty levels and context lengths.
Finally, Appendix §D.2 shows performance on problems categorized by the length of the human-written proof (when available), which we take as a rough proxy of the problem difficulty. The file-tuned model improved on all three difficulty categories. Appendix §D.2 also shows that file-tuning had improved accuracy across context lengths, particularly for problems with longer contexts. Longer contexts may imply more dependencies, suggesting that these problems can benefit more from file-tuning.
Models rely on common symbolic automation.
To demonstrate an additional kind of context-dependence, we perform an additional analysis on Math2001 [15], which contains the Lean code for the book The Mechanics of Proof by Heather Macbeth. [^4] See Appendix B.1 for further details on Math2001. Due to licensing we do not include it in miniCTX. In particular, the textbook code disables powerful automation tactics including simp and nlinarith to promote manual reasoning, akin to traditional textbook exercises. For example, Math 2001 includes numerous arithmetic problems that are trivial with automation tactics (e.g., nlinarith) but are challenging for models to explicitly prove with step-by-step reasoning (e.g., via calc). In Table 5 we evaluate models with the automation disabled, and observe substantial performance drops, confirming a heavy reliance on these automation tactics. An examination of the training corpus further revealed a general dependency on automated tactics within real Lean projects, indicating that our models have learned to rely on these tactics.
5.5 Discussion and future challenges
In addition to general improvements in performance, we comment on some specific open challenges.
Making better use of long-contexts. Our file-tuning method simply truncates contexts to be within a token budget (1024 in our experiments), which can discard useful contextual information. We found gains in providing GPT-4 8,000 tokens of context compared to not providing it context, but its absolute performance was still low. There are several possible strategies that can be explored in future work, including feeding in the entire context, retrieval, or mixtures of the two.
Repository-level context. We focused on evaluating in-file context in this paper. As shown in §D.1, many problems require using context outside of the current file. miniCTX is amenable to studying the cross-file context setting as well. It provides annotations for when a theorem uses context outside of the file, and repository metadata that can be used to support cross-file context.
Challenging proofs. Finally, using context through file tuning did not improve performance on the challenging PFR proofs. Moreover, performance is relatively low (19%) on proofs that had a human-written proof of longer than five lines. Proving these kinds of theorems is an open problem.
Working with constraints. As we showed in Table 5, model performance drops when the proof cannot use standard automation tactics. This can have implications in educational settings such as those in Math 2001, where for instance a more explicit step-by-step proof is desired. Improving performance in this setting of miniCTX is an interesting future direction.
6 Related Work
Formal theorem proving with language models.
GPT-f [9] pioneered the use of language models for theorem proving via next tactic prediction given the current proof state, a technique adopted by many subsequent methods [24, 7, 19, 22, 1, 23]. ReProver [8] conditions each generation on retrieved premises, while Draft-sketch-prove [25] conditions each generation on an informal proof. Baldur [26] fine-tunes a model with 50 lines of the preceding file content as context, but unlike file-tuning trains the model to generate a full proof without proof states. More broadly, machine learning for formal theorem proving is an active research area; see [27, 28] for surveys.
Theorem proving data extraction. Several tools extract training data from interactive theorem provers, including CoqGym [29] for Coq, PISA [24] for Isabelle, LeanStep [7] for Lean 3, and LeanDojo [8] for Lean 3 and 4. Recently, lean-training-data [20] provides tools for extracting proof states and other information using Lean 4 metaprogramming, which we anecdotally found to be easiest to modify and fastest among Lean 4 data extraction tools. Our ntp-toolkit adds 3 new tools on top of this code, along with a pipeline for running on any Lean projects and instruction tuning.
Theorem proving benchmarks. Theorem proving methods are typically evaluated in two settings: (1) standalone competition problems [6] or textbook [30] problems; (2) holding out theorems from a mathematical library that the model is trained on, such as Mathlib for Lean  [7, 22, 8] or the Archive of Formal Proofs for Isabelle [24, 26]. The first does not test the use of context, while the second tests only theorem-level generalization. miniCTX is designed to test the use of context as well as theorem-level, context-level, and project-level generalization across several mathematical domains.
7 Conclusion
We studied the realistic setting of proving theorems that depend on new information and project constraints, and formulated an evaluation framework for testing generalization using real Lean projects. We built miniCTX, and found that the predominant method for training neural theorem provers fails to enable context dependent proving. Our file tuning method provides a strong starting point for the new challenges opened by our investigation into theorem proving with context.
Appendix
Appendix A miniCTX Examples
Here we give some examples of the miniCTX and its sources to illustrate the format of the data and how and why we collect certain theorems.
A.1 Example Entry
An entry in the miniCTX dataset consists of the theorem statement, preceding file contents, and metadata information. For example, given the following theorem s_eq_pow_two in context:
import Mathlib.Data.Real.Basic
# Square function
We define the squaring function s : ℝ‘ to be s x := x * x‘.
def s (x : ℝ) : := x * x
lemma s_eq_pow_two {x : ℝ} : s x = x ^ 2 := by
rw [s, pow_two]
We collect its data into miniCTX, formatted in JSON as follows:
# Preceding file content
[…]
# Theorem statement
[…]
# TheoremName
"theoremName": "s_eq_pow_two",
# Other metadata
"fileCreated": "(gitcommit)",
"theoremCreated": "(gitcommit)",
"file": "(filename)",
"positionMetadata": {
# Line number the theorem is on
"lineInFile": 10,
# Number of tokens before the theorem
"tokenPositionInFile": 152,
# Number of premises (definitions, theorems) before the theorem
"theoremPositionInFile": 1
"dependencyMetadata": {
# Whether the theorem uses definitions or lemmas defined in this file
"inFilePremises": true,
# Whether the theorem uses definitions or lemmas defined in another file in this repository
"repositoryPremises": false
"proofMetadata": {
"hasProof": true,
[…]
"proofType": "tactic",
"proofLengthLines": 2,
"proofLengthTokens": 20
In additional to individual entries, we also record the version (git commit) of the repository.
A.2 PrimeNumberTheoremAnd Example
We collect theorems from the Rectangle.lean file in PrimeNumberTheoremAnd. The following excerpt from Rectangle.lean demonstrates the scenario that often arises in a theorem proving environment where context is critical to producing a proof:
import Mathlib.Analysis.Complex.CauchyIntegral
import Mathlib.Analysis.Complex.Convex
open Complex Set Topology
open scoped Interval
variable {z w : ℂ} {c : ℝ}
[…]
/– A Rectangle has corners z and w‘. -/
def Rectangle (z w : ℂ) : Set := [[z.re, w.re]] ×ℂ [[z.im, w.im]]
namespace Rectangle
lemma symm : Rectangle z w = Rectangle w z := by
simp [Rectangle, uIcc_comm]
lemma symm_re : Rectangle (w.re + z.im * I) (z.re + w.im * I) = Rectangle z w := by
simp [Rectangle, uIcc_comm]
When proving the final lemma symm_re, a model can benefit much from the preceding file contents, which include (1) the existing imports from Mathlib, variable declarations, and open namespaces that provide a syntactic context for this theorem, (2) the new definition Rectangle in the context, which the model has not seen in training, (3) natural language and LaTeX documentation of the file and Rectangle definition, (4) the analogous (in this case identical) proof of the preceding theorem symm. We demonstrate that performance on Rectangle.lean is indeed much higher when preceding file contents are given as context to a model.
For future data added to miniCTX that specifically test the preceding file contents as context, we will ensure it is standalone like Rectangle.lean, i.e. it does not import any other unseen files from the same repository, so the preceding file contents already contain all important information relevant to the proof.
Appendix B Additional datasets
In addition to problems in miniCTX, we also evaluated other datasets that are not included due to copyright reasons.
B.1 Math2001
Math2001  [15] contains the Lean code for the book The Mechanics of Proof by Heather Macbeth, an introductory text on mathematical theorem proving with accompanying Lean code. Each chapter of The Mechanics of Proof covers an introductory topic and walks through how to write the associated mathematics in Lean, along with exercises. The topics include proofs by calculation, proofs with structure, parity and divisibility, logic, induction, number theory, functions, sets, and relations. A unique aspect of Math2001 is that it disables common Lean automation for pedagogical purposes. For example, a student must write out an equality proof in detail, with each step justified. It also defines new tactics and definitions separate from the common Lean libraries. Typically a file in the textbook will show examples of such proofs, followed by exercises for a student to complete. We can view this as a form of contextual adaptation: a model must prove the theorem according to the constraints of the textbook. Math2001 has 41 files that include examples and exercises. We selected 1 to 2 theorems from each file (depending on the length of the file), for a total of 50 theorems. Of these, 31 have no proof in the Math2001 repository, hence testing theorem-level generalization.
Context-aware models surpass state-based models
Table 4 shows the performance comparison of different models. Both the GPT-4o model, which includes context in the input, and the file-tuned model perform significantly better than the other models. This demonstrates the importance of context information in context-dependent textbook-style problems.
Models rely on common symbolic automation.
The Math2001 split originally disables powerful automation tactics including simp and nlinarith to promote manual reasoning, akin to traditional textbook exercises. In Table 5 we evaluate models with the automation disabled, and observe substantial performance drops, confirming a heavy reliance of current models on these automation tactics. An examination of the training corpus further revealed a general dependency on automated tactics within real Lean projects, indicating that our models have learned to rely on these tactics.
Models Math2001
GPT-4o (full proof) 11.76%
GPT-4o (+ context) 43.13%
State-tactic prompting 31.37%
State-tactic tuning 27.45%
File tuning 41.18%
Table 4: Performance comparison of different models on Math2001.
Automation File (%) State-tactic (%)
Enabled 41.18 11.76
Disabled 27.45 7.84
Table 5: Performance on the Math2001 split with and without access to standard automation.
Appendix C ntp-toolkit and file-tuning details
C.1 Data extraction
ntp-toolkit contains a general-purpose data extraction tool that extracts examples from an arbitrary Lean 4 repository and formats them into examples for language-model fine-tuning. The tool is implemented in Lean based on Kim Morrison’s lean-training-data.
Specifically, ntp-toolkit takes in a configuration file with one or more Lean repositories specified. Each repository is transformed into next-tactic and full proof examples stored in JSON Lines files. The next-tactic data is suitable for making file-tuning examples of the form (context, state, next-tactic):
"state": # tactic state ,
"nextTactic" : # pretty-printed next tactic,
"srcUpToTactic" : # source code in the file up to the tactic invocation,
"decl": # declaration without proof (e.g., statement of a theorem),
"declUpToTactic" : # source code in the declaration up to the tactic invocation,
"declId": # unique identifier of the declaration
The full proof data is suitable for making evaluation examples of the form (context, theorem, proof):
"srcUpToDecl": # source code in the file up to the declaration,
"decl": # declaration without proof (e.g., statement of a theorem),
"declId": # unique identifier of the declaration,
"proof": # proof
Full proof data is also suitable for training a model to directly generate a full proof, and ntp-toolkit also provides Lean source with proof states interleaved, both of which we do not explore in this work.
C.2 Input-output formatting.
Below we show the inputs and outputs for file-tuning and state-tactic tuning. In the paper we refer to the natural language description at the beginning of the input as an “instruction”, and refer to a set of inputs and outputs as described below as “instruction-tuning data”.
C.2.1 File tuning.
Given an example containing a state, next-tactic, and preceding file contents (srcUpToTactic), the data is formatted as:
Input:
/- You are proving a theorem in Lean 4.
You are given the following information:
- The file contents up to the current tactic, inside [CTX]...[/CTX]
- The current proof state, inside [STATE]...[/STATE]
Your task is to generate the next tactic in the proof.
Put the next tactic inside [TAC]...[/TAC]
[CTX]
{srcUpToTactic}
[/CTX]
[STATE]
{state}
[/STATE]
[TAC]
Output:
{nextTactic}
[/TAC]
C.2.2 State-tactic tuning.
Given an example containing a state and next-tactic, the data is formatted as:
Input:
/- You are proving a theorem in Lean 4.
You are given the following information:
- The current proof state, inside [STATE]...[/STATE]
Your task is to generate the next tactic in the proof.
Put the next tactic inside [TAC]...[/TAC]
[STATE]
{state}
[/STATE]
[TAC]
Output:
{nextTactic}
[/TAC]
C.2.3 GPT4-o prompt
For full proof generation task with only theorem statement, we use the following prompt:
Your task is to generate complete proofs for problems stated in Lean4. You may use any tactics available in Mathlib, but no additional context, definitions, or theorems from the problem’s file will be provided. Focus on crafting proofs using general knowledge and techniques applicable in Lean4. Here are some examples: lemma deriv_scale {f : CS (n + 1) E} : (f.scale R).deriv = R⁻¹ f.deriv.scale R := by ext v ; by_cases hR : R = 0 <;> simp [hR, scale] · simp [deriv, smul] ; exact deriv_const _ _ · exact ((f.hasDerivAt (R⁻¹ v)).scomp v (by simpa using (hasDerivAt_id v).const_smul R⁻¹)).deriv theorem mul_dvd_mul_left (a : α) (h : b | c) : a * b | a * c := by obtain d, rfl := h use d rw [mul_assoc] /- Now here is your exercise. There is no need to restate the problem. If needed, think through the proof using comments. -/ {theorem statement}
For full proof generation task with additional infile context, we use the following prompt:
Your task is to generate complete proofs for problems stated in Lean4. For each problem, you will be provided with the context from the file in which the theorem is stated. This context includes useful external libraries, along with important definitions and theorems that are relevant to the proof. You are encouraged to use any tactics, definitions, lemmas, or theorems defined within this context to construct your proof. Please pay careful attention to indentation and formatting to ensure that the proof adheres to Lean4 syntax standards. Here are some examples: #Context: import Mathlib.Analysis.Calculus.Deriv.Support import Mathlib.Analysis.Distribution.SchwartzSpace import Mathlib.Order.Filter.ZeroAndBoundedAtFilter open Real Complex MeasureTheory Filter Topology BoundedContinuousFunction SchwartzMap BigOperators variable {{E : Type*}} [NormedAddCommGroup E] [NormedSpace E] {{n : ℕ}} @[ext] structure CS (n : ℕ) (E : Type*) [NormedAddCommGroup E] [NormedSpace E] where toFun : E h1 : ContDiff n toFun h2 : HasCompactSupport toFun noncomputable def scale (g : CS n E) (R : ℝ) : CS n E := by by_cases h : R = 0 · exact 0, contDiff_const, by simp [HasCompactSupport, tsupport] · refine fun x => funscale g R x, ?_, ?_ · exact g.h1.comp (contDiff_const.smul contDiff_id) · exact g.h2.comp_smul (inv_ne_zero h) /- Truncated -/ /- Now here is your exercise. There is no need to restate the problem. If needed, think through the proof using comments. -/ #Context: {} #Problem: {} {theorem statement}
Appendix D Additional results and analysis
D.1 Label distribution
Figure 4: Percentage of different dependencies in miniCTX
D.2 Performance by proof length and context length.
Figure 5: Performance of two methods on different difficulty levels and context lengths
D.3 Example of modifying proofs from context
The File-tuned model is able to utilize the proofs in the context. Here is an example of the model making minimal modification to the proofs from the context:
Input:
theorem Set.left_not_mem_uIoo {a b : ℝ} : a Set.uIoo a b := fun h1, h2 (left_lt_sup.mp h2) (le_of_not_le (inf_lt_left.mp h1)) theorem Set.right_not_mem_uIoo {a b : ℝ} : b Set.uIoo a b := by
Output:
rintro h1, h2 exact (right_lt_sup.mp h2) (le_of_not_le (inf_lt_right.mp h1))
D.4 Example of using theorems from context
The File-tuned model is able to utilize the stated theorems in the context. Here is an example of the model using the previously defined theorem in the proof:
Input:
lemma RectSubRect {x x x x y y y y : ℝ} (x_le_x : x x₁) (x_le_x : x x₂) (x_le_x : x x₃) (y_le_y : y y₁) (y_le_y : y y₂) (y_le_y : y y₃) : Rectangle (x + y * I) (x + y * I) Rectangle (x + y * I) (x + y * I) := by rw [rect_subset_iff, mem_Rect, mem_Rect] refine ⟨⟨?_, ?_, ?_, ?_⟩, ?_, ?_, ?_, ?_ all_goals simpa using by linarith lemma RectSubRect {z z z z : ℂ} (x_le_x : z₀.re z₁.re) (x_le_x : z₁.re z₂.re) (x_le_x : z₂.re z₃.re) (y_le_y : z₀.im z₁.im) (y_le_y : z₁.im z₂.im) (y_le_y : z₂.im z₃.im) : Rectangle z z Rectangle z z := by
Output:
simpa using RectSubRect x_le_x x_le_x x_le_x y_le_y y_le_y y_le_y
D.5 Example of using unseen definitions
The State-tactic-tuned model is able to utilize the even unseen definitions that appear in the proof state. In the following example Set.uIoo is a newly defined definition, which is never seen for state-tactic tuned model:
Input:
theorem uIoo_of_le : Type*} [Lattice α] {a b : α} (h : a b) : Set.uIoo a b = Ioo a b := by [STATE] 1 goal zw: c: α: Type u_1 inst: Lattice α ab: α h: a b uIoo a b = Ioo a b [/STATE]
Output:
simp [uIoo, h]
Appendix E Dataset Hosting, Licensing, and Maintenance
miniCTX is released on HuggingFace: l3lab/miniCTX [^5] Croissant metadata record: https://huggingface.co/api/datasets/l3lab/miniCTX/croissant., distributed under the Apache 2.0 license. Data extraction tool ntp-toolkit is released on GitHub: cmu-l3/ntp-toolkit, under the MIT license. We note that the underlying data for the individual splits of miniCTX are also released under the Apache 2.0 license. We include the licensing information in the dataset repository. We plan to regularly update and maintain the dataset to include examples from new projects.
Appendix F Annotation Guideline
With the NTP-TOOLKIT, users can extract and annotate new theorems from any valid Lean project.
F.1 Preliminary
The evaluation code relies heavily on the Lean REPL, which operates within the project environment. Therefore, it is essential that the project builds without any errors. Additionally, the version of Lean used in the project should match the version supported by the REPL. While the Lean REPL supports versions $\geq$ 4.3.0, for the best experience, we recommend evaluating projects that use Lean versions between 4.5.0 and 4.7.0.
F.2 Using the NTP-TOOLKIT
The NTP-TOOLKIT is designed to facilitate the extraction and annotation of theorem proving data from Lean projects. To use the toolkit for annotating theorems, follow these steps:
  1. 1.
    Installation: Clone the NTP-TOOLKIT repository from GitHub to your local machine. Ensure that you have the required dependencies installed, as listed in the repository’s README file.
  2. 2.
    Configuration: Configure the toolkit to point to the directory of your Lean project. Make sure that your project is using a compatible version of Lean as per the guidelines in the previous section.
  3. 3.
    Data Extraction: Run the data extraction script provided by the toolkit. Specify the full_proof_training_data option to focus the extraction on datasets suitable for training models with full proofs.
For detailed commands and additional options, please refer to the README file in the NTP-TOOLKIT repository.
F.3 Evaluation
We provide a comprehensive evaluation pipeline in the miniCTX-eval repository, supporting both tactic-prediction and full-proof generation tasks. Users should place the extracted JSONL file from the NTP-TOOLKIT into the data folder. To run an evaluation task, execute the task script by specifying the dataset path, the corresponding project path, and the path to the Lean REPL. This setup ensures that the evaluation is conducted within the correct environment and with the necessary data inputs.